OpenStudy (anonymous):

In the class elections, Connie received 53 more votes than Carl. In all, 221 votes were cast for the two candidates. How many votes did each receive. One point for the equation, one point for the solution.

OpenStudy (anonymous):

let connie = g and carl = b... g+b = 221 g = b +53. those are the equations you should be able to solve now, if you cant let me know ill work you thru it

OpenStudy (anonymous):

i don't' get how to do it

OpenStudy (anonymous):

can you help me. please and thank you.

OpenStudy (anonymous):

since g = b + 53 plug g into the other equation to have on variable to solve... so, (b + 53) + b = 221... 2g+53=221,... subtract 53 from both side,... 2g = 168,... dividen both side by 2 to show that g=84. so karl gor 84 votes. plug in 84 for g on the equation b+g=221. .. so 84 + G = 221. solving for that we can see connie ( or g) got 137 votes. sorry i know this is sloppy but im at work so its hard to do this ahahh but i hope i helped.!

OpenStudy (anonymous):

sorry i noticed that after i plugged in for b + 53 i switched the variables but you get the idea i hope!

OpenStudy (anonymous):

yea thank you